OpenStudy (anonymous):

Evaluate loga24 using the fact that loga2 = 0.4307 and loga3 = 0.6826

OpenStudy (anonymous):

use the fact that \(24=2^3\times 3\)

OpenStudy (anonymous):

so you want \[\log(2^3\times 3)\] then use the properties of the log to expand this as a sum and product of logs

OpenStudy (anonymous):

step one is \[\log(2^3\times 3)=\log(2^3)+\log(3)\] since \[\log(AB)=\log(A)+\log(B)\] always

OpenStudy (anonymous):

Thank you.